OpenStudy (anonymous):

Find the last 3 digits of 7^12341. I know that the last 3 digits are 7^12341 (mod10^3) but I can't see how to figure that out

ganeshie8 (ganeshie8):

As a start : find \(\large \phi(10^3)\)

ganeshie8 (ganeshie8):

next use the euler theorem and reduce : \[\large a^{\phi(n)} \equiv 1 \pmod n\]

ganeshie8 (ganeshie8):

Notice that \(\large \gcd(7,10^3)=1\) allowing you to use euler thm

OpenStudy (anonymous):

oh so we do need Euler's? I read a bit about it but couldn't remember my prof going over it. Was sick one day though....okay.

ganeshie8 (ganeshie8):

you familiar with Fermat little theorem ?

OpenStudy (dan815):

O_O

OpenStudy (anonymous):

Not really, most advanced I can recall is Hensella's Lemma

ganeshie8 (ganeshie8):

ahh then forget about euler and fermat, you can use Hensel's lemma along with chinese remainder theorem

OpenStudy (anonymous):

oh? Could you explain how?

ganeshie8 (ganeshie8):

split it into two congruences : 7^12341 (mod 2^3) 7^12341 (mod 5^3)

ganeshie8 (ganeshie8):

what does Hensel's lemma tell you ?

OpenStudy (anonymous):

tells you how to find a solution to the polynomial P(x) such that it is congruent to 0 (modp^n) where p is prime

OpenStudy (anonymous):

midterm on thursday, he wanted this section on it so he rushed through both chinese remainder theorem and Hensel's lemma

ganeshie8 (ganeshie8):

Alright, so you want to solve below first 7^12341 (mod 2) 7^12341 (mod 5) is it ?

OpenStudy (anonymous):

since 7 is odd, 7^12341 (mod 2) will be 1 for mod 5....I believe I have that 7^12341 (mod 5) is congruent with 7 (mod 5) right? which is just 2 (mod 5)

OpenStudy (anonymous):

or did I get your question confused....

ganeshie8 (ganeshie8):

yes how do you work mod 2^3 having known mod 2 ? let me confess, i haven't learned hensel's lemma in my number theory..

ganeshie8 (ganeshie8):

actually 7^12341 (mod 2^3) is straightforward as 7 = -1 mod 8 : 7^12341 = (-1)^12341 = 7 mod 2^3

OpenStudy (anonymous):

thats the thing though, Hensel's lemma only applies to solving polynomials that are set to be congruent to 0 modulo a prime to a power

ganeshie8 (ganeshie8):

if we know how to find 7^12341(mod 5^3) , we can use chinese remainder thm and conclude

OpenStudy (anonymous):

I see you got the right answer there, but I don't know how you knew to use that it was 7 mod 8 and not say 5 or 3 or 1

ganeshie8 (ganeshie8):

(-1)^(odd number) = -1 = 7 (mod 8)

OpenStudy (anonymous):

oh jeeze, had nothing to do with the modulo 2 thing okay

ganeshie8 (ganeshie8):

yeah we need to figure out how to reduce something mod prime power... i have been skipping this topic from months.. :o

OpenStudy (anonymous):

lol, let me reread CRT, I suspect thats what is needed.

ganeshie8 (ganeshie8):

we can apply CRT only after reducing 7^12341 (mod 5^3)

OpenStudy (anonymous):

isn't there a unnamed theorem about a^n = a (mod b) if a and b are relatively prime?

ganeshie8 (ganeshie8):

thats the euler theorem which we don't want to use : \(a^{\phi(n)} \equiv 1 \pmod n\)

OpenStudy (anonymous):

Alright, read a bit ahead trying to solve this. Probably remembering it from there.

OpenStudy (anonymous):

Might just end up skipping this and asking a friend in my class tomorrow. If I do I will post how to do it on here tomorrow for those interested.

ganeshie8 (ganeshie8):

Okay good luck with rest of the preparation :) here is a way to work it using euler theorem : \[\large \phi(10^3) = \phi(2^3)\phi(5^3) = (4)(100) = 400\] \[\large 7^{12341} \equiv 7^{400\times 30 + 341}\equiv 7^{341} \pmod {1000}\]

ganeshie8 (ganeshie8):

you can reduce 7^341 as follows : 7^2 = 49 7^4 = 49*49 = 401 7^8 = 401*401 = 801 7^16 = 801*801 = 601 7^32 = 601*601 = 201 ... 7^256 = 601 7^341 = 7^(256 + 64 + 16 + 4 + 1 ) = 601*401*601*401*7 (mod 1000) = 007

ganeshie8 (ganeshie8):

which is indeed a special number to have as last 3 digits, but im sure there must be more efficient way... do let me know when you find it :)

OpenStudy (anonymous):

The way my professor intended this question to be completed was through CRT and looking at (mod8) and (mod 125). mod 8 has an easy pattern to see 7^1 is congruent to 7 which is congruent to -1 (mod 8) 7^2 is congruent to 1 (mod 8) Therefore 7^12341 follow the pattern of odd powers of 7 and is congruent to 7 (mod 8) (mod 125) is tougher, and you need to do 10 powers of 7 to get the pattern. I have a test tomorrow and cannot spend the time to write these guys out but it suffices to say that there is a pattern that appears at 7^10 as 7^10 is congruent to -1. Using this pattern we can determine that 7^12341 is congruent to 7 (mod 125) From the CRT we get that 7^12341 is congruent to 7 (mod 1000) thus the last 3 digits of 7^12341 is 007. @ganeshie8 sorry it was a bit late, been busy studying for the test.

ganeshie8 (ganeshie8):

basically you're looking for what power of 7 gives you a 1 or -1 and using it to reduce. Nice :)